OpenStudy (anonymous):

A customer from Cavallaro's Fruit Stand picks a sample of 3 oranges at random from a crate containing 60 oranges, of which 4 are rotten. What is the probability that the sample contains 1 or more rotten oranges? (Round your answer to three decimal places.)

OpenStudy (anonymous):

So if it has at least a rotten egg in those 5, we can easily find the probability of getting all fresh eggs P(1st egg fresh) * P(2nd) * P(3rd) * P(4th) = (60/65) * (59/64) * (58/63) * (57/62) = (60*59*58*57) / (65*64*63*62) = 0.72024548 1 - 0.72024548 = 0.27975452 = approx. 0.280 The probability of getting at least rotten in 4 oranges is 0.28 or 28%

OpenStudy (anonymous):

I got this answer too when I tried it myself, but it comes up incorrect in Webassign. I put 0.280 @jrc5101520

OpenStudy (anonymous):

what is webassign

OpenStudy (anonymous):

It's the website I have to do my homework through for my class

OpenStudy (anonymous):

that is wearied mabay you should talk to your teacher

OpenStudy (anonymous):

= 1 - 0.7757 ---- = 0.2243

OpenStudy (anonymous):

@rusalka